# RestoreVAR: Visual Autoregressive Generation for All-in-One Image Restoration

### v2

@@ Abstract
The use of latent diffusion models (LDMs) such as Stable Diffusion has significantly improved the perceptual quality of All-in-One image Restoration (AiOR) methods, while also enhancing their generalization capabilities. However, these LDM-based frameworks suffer from slow inference due to their iterative denoising process, rendering them impractical for time-sensitive applications. Visual autoregressive modeling (VAR), a recently introduced approach for image generation, performs scale-space autoregression and achieves comparable performance to that of state-of-the-art diffusion transformers with drastically reduced computational costs. Moreover, our analysis reveals that coarse scales in VAR primarily capture degradations while finer scales encode scene detail, simplifying the restoration process. Motivated by this, we propose RestoreVAR, a novel VAR-based generative approach for AiOR that significantly outperforms LDM-based models in restoration performance while achieving over $\mathbf{10\times}$ faster inference. To optimally exploit the advantages of VAR for AiOR, we propose architectural modifications and improvements, including intricately designed cross-attention mechanisms and a latent-space refinement module, tailored for the AiOR task. Extensive experiments show that RestoreVAR achieves state-of-the-art performance among generative AiOR methods, while also exhibiting strong generalization capabilities.
@@ 1 Introduction
Image restoration is a complex inverse problem that aims to recover clean images from degradations, such as haze, rain, snow, blur, and low-light conditions. Recently, the paradigm of All-in-One image Restoration (AiOR) has emerged, where a single network is trained to handle multiple degradation types. Existing AiOR methods can be broadly categorized into non-generative and generative approaches. Non-generative models such as AirNet (Li et al., 2022), PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan & Patel, 2024), and AdaIR (Cui et al., 2024), deterministically map degraded images to their clean counterparts. While these methods offer fast inference and reliable pixel-level restoration performance, they often fail to generalize to diverse degradations encountered in real-world scenarios. To overcome this challenge, recent works have adopted generative models that aim to capture the distribution of clean images and produce more perceptually realistic outputs. Early works (Chen et al., 2022; Kupyn et al., 2018) based on GANs (Goodfellow et al., 2020) attempted this through adversarial learning, but suffered from mode collapse and unstable training. To improve fidelity and training stability, DiffUIR (Zheng et al., 2024) and DA-CLIP (Luo et al., 2023) employed pixel-space diffusion models (Ho et al., 2020). However, their high computational cost makes large-scale pretraining infeasible, limiting their ability to learn strong generative priors. In contrast, recent methods such as Diff-Plugin(Liu et al., 2024), AutoDIR(Jiang et al., 2023), and PixWizard (Lin et al., 2024) leverage latent diffusion models (LDMs), such as Stable Diffusion (Rombach et al., 2022). By operating in a latent space, LDMs significantly reduce computational costs, enabling large-scale pretraining which equips them with strong generative priors of natural images. These priors allow LDM-based AiOR methods to deliver perceptually realistic restoration and improved generalization to real-world degradations.
@@
Autoregressive models have driven rapid advances in natural language processing through large language models (LLMs) such as GPT-3 (Radford et al., 2019b) and LLaMA (Touvron et al., 2023). These models generate outputs by predicting the next token, conditioned on previously generated tokens. Recently, Visual AutoRegressive (VAR) Modeling (Tian et al., 2024) introduced scale-space autoregression for image generation, performing next-scale prediction in the latent space of a multi-scale vector-quantized VAE (VQVAE). VAR achieves performance comparable to state-of-the-art diffusion models such as DiT-XL/2 (Peebles & Xie, 2023), while operating $45\times$ faster. Despite its success in generative tasks, the application of VAR to low-level vision tasks such as image restoration remains largely unexplored. To the best of our knowledge, only two prior works—VarSR (Qu et al., 2025) and Varformer (Wang & Zhao, 2024)—have used VAR for image restoration. VarSR focused exclusively on the super-resolution task, while Varformer utilized intermediate VAR features to guide a separate non-generative network for AiOR. In contrast, our approach is generative and fully exploits the strong priors of the pretrained VAR model by training it directly for the AiOR task. Our analysis in Sec. 3.2 also reveals that the scale-space decomposition of VAR captures degradations predominantly in coarse scales and scene-level details in fine scales, making it well-suited for AiOR.
@@ 2.1 Image restoration
Early restoration models primarily addressed specific degradations  (He et al., 2009; Zhang et al., 2020; Wang et al., 2019; Yasarla & Patel, 2019; Zhang et al., 2021a; Nah et al., 2017). Other methods such as Restormer (Zamir et al., 2022), MPRNet (Zamir et al., 2021) and SwinIR (Liang et al., 2021) introduced architectures for any single restoration task. However, they are restricted to handle one degradation at a time, making them ineffective for multiple degradations. All-in-One image Restoration (AiOR) methods aim to tackle multiple corruptions with a single model. Early approaches include non-generative models such as All-in-one (Li et al., 2020) and Transweather (Valanarasu et al., 2022). PromptIR (Potlapalli et al., 2024) used learnable prompts while AWRaCLe (Rajagopalan & Patel, 2024) utilized visual in-context learning to extract degradation characteristics. Other approaches such as InstructIR (Conde et al., 2025) adopted textual guidance, and DCPT (Hu et al., 2025) proposed a novel pre-training strategy for AiOR. DFPIR (Tian et al., 2025) proposed a feature perturbation strategy for AiOR. Recent AiOR methods have adopted diffusion models. Pixel-space diffusion models (PSDMs) such as DA-CLIP (Luo et al., 2023) and DiffUIR (Zheng et al., 2024) demonstrated improved AiOR performance but lacked robust generative priors. Recent methods have utilized the strong priors of LDMs for AiOR. Diff-Plugin (Liu et al., 2024) adopts task plugins to guide an LDM for AiOR. AutoDIR (Jiang et al., 2023) automatically detects and restores degradations using an LDM. PixWizard (Lin et al., 2024) is a multi-task SD-XL (Podell et al., 2023) based model capable of performing AiOR among other tasks. However, LDM-based approaches are slow at inference time—a limitation we aim to overcome using visual autoregressive modeling (VAR).
@@ 2.2 Autoregressive models in vision
Recent works (Van Den Oord et al., 2016; Tian et al., 2024) have extended autoregressive (AR) models to vision and can be categorized as pixel‐space AR (Van Den Oord et al., 2016; Van den Oord et al., 2016; Chen et al., 2018), token‐based AR (Van Den Oord et al., 2017; Yu et al., 2023; Ramesh et al., 2021) and scale‐space AR (Tian et al., 2024; Ren et al., 2024; Guo et al., 2025). Pixel-space AR predicts raw pixels one by one in raster order, as in PixelRNN (Van Den Oord et al., 2016) and PixelCNN++ (Salimans et al., 2017), but is very slow at high resolutions. Token-based AR compresses images into discrete latent codes via vector quantization (e.g., VQ-VAE (Van Den Oord et al., 2017), VQGAN (Esser et al., 2021)) and then models code sequences with transformers (e.g. ImageGPT (Chen et al., 2020)). This trades-off codebook size and transformer capacity against tractability for high-resolution generation. Scale-space AR, as introduced in VAR (Tian et al., 2024), generates latents from coarse to fine scales and matches the quality of Diffusion Transformers (Peebles & Xie, 2023) at a fraction of the inference cost. HART (Tang et al., 2024) scales VAR to higher resolution and uses a MLP-based diffusion refiner to convert discrete VAR latents into continuous representations. Despite VAR’s success in generative tasks, it remains underexplored for image restoration with only two prior works-VarSR (Qu et al., 2025) and Varformer (Wang & Zhao, 2024). VarSR addressed super-resolution, while Varformer used VAR’s features to guide a non-generative AiOR model. In contrast, RestoreVAR is a generative model which directly trains VAR for AiOR.
@@ 3.1 Preliminaries: Visual Autoregressive Modelling
Visual Autoregressive Modelling, or VAR, is a novel autoregressive class-conditioned image generation method which uses a GPT-2 (Radford et al., 2019a) style decoder-only transformer architecture for next-scale prediction. The VAR transformer operates in the latent space of a multi-scale VQVAE which uses $K$ scales. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the VQVAE encoder outputs a latent representation $f_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. Hereafter, we will refer to $f_{\text{cont}}$ as the continuous latent, and the latent obtained after quantization as discrete latent. Instead of directly quantizing $f_{\text{cont}}$, a multi-scale residual quantization using a shared codebook across $K$ spatial scales is performed. First, the residual and accumulated quantized (or discrete) reconstruction of $f_{\text{cont}}$ are initialized as $f_{\text{res}}^{(0)}:=f_{\text{cont}}$ and $f_{\text{quant}}^{(0)}:=0$, respectively. At each scale $k=1,\ldots,K$, an index map $r_{k}\in\mathbb{Z}^{H_{k}\times W_{k}}$ is obtained by quantizing the downsampled residual feature:
@@ 3.3 RestoreVAR
Degraded GT++coarse GT++fine GT
Haze
Refer to caption Refer to caption Refer to caption Refer to caption
Snow
Refer to caption Refer to caption Refer to caption Refer to caption
Rain
Refer to caption Refer to caption Refer to caption Refer to caption
Low-light
Refer to caption Refer to caption Refer to caption Refer to caption
Blur
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 3.3.2 Detail-Preserving Restoration
$\mathcal{L}_{\text{dec}}=\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{percep}}+\lambda_{4}\mathcal{L}_{\text{adv}},\vskip-4.0pt$
@@
$\hat{f}_{\text{cont}}=f^{\text{pred}}_{\text{quant}}+\text{LRM}(f^{\text{pred}}_{\text{quant}},z),\vskip-4.0pt$
@@ 4.3 Comparisons
We compare RestoreVAR with state-of-the-art generative and non-generative methods for AiOR. For non-generative approaches, we include PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan & Patel, 2024), DCPT (Hu et al., 2025) and DFPIR (Tian et al., 2025). Among generative methods, we compare with the LDM-based approaches Diff-Plugin (Liu et al., 2024), AutoDIR (Jiang et al., 2023) and PixWizard (Lin et al., 2024). To ensure a fair comparison, we retrained PromptIR and AWRaCLe, as their official checkpoints were not trained for most of our AiOR tasks. All other methods were evaluated using their publicly released checkpoints. For AutoDIR, we report results without the structure correction module, as this module functions as an independent, non-generative restoration network (more details in supplementary). The results reported for PixWizard were obtained using its publicly released checkpoint. We do not compare with task-specific restoration models, as RestoreVAR is proposed for the AiOR setting.
Table 4.2 presents PSNR, SSIM and LPIPS scores on the RESIDE, Snow100k, Rain13K, LOLv1 and GoPro datasets. RestoreVAR surpasses LDM-based AiOR methods at a fraction of their computational cost (inference time (s) per image)—Diff-Plugin:$2.04$s, AutoDIR: $8.477$s, PixWizard: $8.247$s and RestoreVAR: $0.201$s, highlighting the efficacy of our framework. More detailed complexity comparisons are given in the supplementary along with a derivation showing that the time complexity of VAR with maximum latent resolution $n\times n$ is $\mathcal{O}(\log n)$ lower than an LDM operating at the same latent resolution. Qualitative comparisons with LDM-based methods in Fig. 5 further illustrate that RestoreVAR produces restored images of high quality while better preserving fine details. Visual results for the Snow100k and LOLv1 datasets are provided in the supplementary. While non-generative methods achieve better scores, it is important to recognize that the performance of RestoreVAR is inherently influenced by the quality of the VAE decoder; a limitation shared by all latent generative approaches. Despite this constraint, RestoreVAR narrows the gap with non-generative methods while maintaining the benefits of a generative framework, i.e., perceptually realistic results and strong generalization capabilities. To demonstrate these strengths, we evaluate generalization using no-reference image quality metrics (following prior works (Liu et al., 2024; Jiang et al., 2023; Rajagopalan & Patel, 2024)), and assess perceptual realism through a user study.
Table 2: Quantitative comparisons of RestoreVAR against state-of-the-art non-generative approaches on real-world, unseen and mixed degradations. The best result is indicated in bold.
@@
For testing generalization, we report MUSIQ (Ke et al., 2021) and CLIPIQA (Wang et al., 2023) scores in Table 2 on the real-world, unseen and mixed degradation datasets discussed in Sec. 4.2.
@@
RestoreVAR achieves higher scores than non-generative models (on average), indicating better robustness under these degradations. Qualitative results for this experiment are shown in Fig. 6, where RestoreVAR consistently outperforms non-generative approaches. Due to space constraints, qualitative comparisons with PromptIR and visual results for LOLBlur are given in the supplementary. To further evaluate perceptual quality, we conducted a user study in which participants rated outputs from non-generative models, AutoDIR (LDM-based) and RestoreVAR, for $50$ real-world scenes. We received $36$ responses with each participant scoring outputs based on scene consistency, restoration quality, and overall appeal on a 5-point scale. Table 3 shows that RestoreVAR received the highest average ratings (across all three criteria), highlighting its ability to produce images that align closely with human preferences.
These results highlight the effectiveness of RestoreVAR for AiOR. We discuss limitations of RestoreVAR in the supplementary.
4.4 Ablations
Discriminator for VAE fine-tuning. As described in Sec. 3.3.2, we fine-tune the VAE decoder on continuous latents using a combination of pixel-level loss and an adversarial loss. To analyze the impact of the discriminator, we compare the reconstructions of VAE decoders fine-tuned with and without the adversarial loss. As shown in Fig. 8, removing the discriminator leads to blurrier reconstruction while including it yields sharper and perceptually better looking outputs.

### v3

@@ Abstract
The use of latent diffusion models (LDMs) such as Stable Diffusion has significantly improved the perceptual quality of All-in-One image Restoration (AiOR) methods, while also enhancing their generalization capabilities. However, these LDM-based frameworks suffer from slow inference due to their iterative denoising process, rendering them impractical for time-sensitive applications. Visual autoregressive modeling (VAR), a recently introduced approach for image generation, performs scale-space autoregression and achieves comparable performance to that of state-of-the-art diffusion transformers with drastically reduced computational costs. Moreover, our analysis reveals that coarse scales in VAR primarily capture degradations while finer scales encode scene detail, simplifying the restoration process. Motivated by this, we propose RestoreVAR, a novel VAR-based generative approach for AiOR that significantly outperforms LDM-based models in restoration performance while achieving over $\mathbf{10\times}$ faster inference. To optimally exploit the advantages of VAR for AiOR, we propose architectural modifications and improvements, including intricately designed cross-attention mechanisms and a latent-space refinement module, tailored for the AiOR task. Extensive experiments show that RestoreVAR achieves state-of-the-art performance among generative AiOR methods, while also exhibiting strong generalization capabilities. Project Page: https://sudraj2002.github.io/restorevarpage/
@@ 1 Introduction
Image restoration is a complex inverse problem that aims to recover clean images from degradations, such as haze, rain, snow, blur, and low-light conditions. Recently, the paradigm of All-in-One image Restoration (AiOR) has emerged, where a single network is trained to handle multiple degradation types. Existing AiOR methods can be broadly categorized into non-generative and generative approaches. Non-generative models such as AirNet (Li et al., 2022), PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan and Patel, 2024), and AdaIR (Cui et al., 2024), deterministically map degraded images to their clean counterparts. While these methods offer fast inference and reliable pixel-level restoration performance, they often fail to generalize to diverse degradations encountered in real-world scenarios. To overcome this challenge, recent works have adopted generative models that aim to capture the distribution of clean images and produce more perceptually realistic outputs. Early works (Chen et al., 2022; Kupyn et al., 2018) based on GANs (Goodfellow et al., 2020) attempted this through adversarial learning, but suffered from mode collapse and unstable training. To improve fidelity and training stability, DiffUIR (Zheng et al., 2024) and DA-CLIP (Luo et al., 2023) employed pixel-space diffusion models (Ho et al., 2020). However, their high computational cost makes large-scale pretraining infeasible, limiting their ability to learn strong generative priors. In contrast, recent methods such as Diff-Plugin(Liu et al., 2024), AutoDIR(Jiang et al., 2023), and PixWizard (Lin et al., 2024) leverage latent diffusion models (LDMs), such as Stable Diffusion (Rombach et al., 2022). By operating in a latent space, LDMs significantly reduce computational costs, enabling large-scale pretraining which equips them with strong generative priors of natural images. These priors allow LDM-based AiOR methods to deliver perceptually realistic restoration and improved generalization to real-world degradations.
@@
Autoregressive models have driven rapid advances in natural language processing through large language models (LLMs) such as GPT-3 (Radford et al., 2019b) and LLaMA (Touvron et al., 2023). These models generate outputs by predicting the next token, conditioned on previously generated tokens. Recently, Visual AutoRegressive (VAR) Modeling (Tian et al., 2024) introduced scale-space autoregression for image generation, performing next-scale prediction in the latent space of a multi-scale vector-quantized VAE (VQVAE). VAR achieves performance comparable to state-of-the-art diffusion models such as DiT-XL/2 (Peebles and Xie, 2023), while operating $45\times$ faster. Despite its success in generative tasks, the application of VAR to low-level vision tasks such as image restoration remains largely unexplored. To the best of our knowledge, only two prior works: VarSR (Qu et al., 2025) and Varformer (Wang and Zhao, 2024), have used VAR for image restoration. VarSR focused exclusively on the super-resolution task, while Varformer utilized intermediate VAR features to guide a separate non-generative network for AiOR. In contrast, our approach is generative and fully exploits the strong priors of the pretrained VAR model by training it directly for the AiOR task. Our analysis in Sec. 3.2 also reveals that the scale-space decomposition of VAR captures degradations predominantly in coarse scales and scene-level details in fine scales, making it well-suited for AiOR.
@@ 2.1 Image restoration
Early restoration models primarily addressed specific degradations  (He et al., 2009; Zhang et al., 2020; Wang et al., 2019; Yasarla and Patel, 2019; Zhang et al., 2021a; Nah et al., 2017). Other methods such as Restormer (Zamir et al., 2022), MPRNet (Zamir et al., 2021) and SwinIR (Liang et al., 2021) introduced architectures for any single restoration task. However, they are restricted to handle one degradation at a time, making them ineffective for multiple degradations. All-in-One image Restoration (AiOR) methods aim to tackle multiple corruptions with a single model. Early approaches include non-generative models such as All-in-one (Li et al., 2020) and Transweather (Valanarasu et al., 2022). PromptIR (Potlapalli et al., 2024) used learnable prompts while AWRaCLe (Rajagopalan and Patel, 2024) utilized visual in-context learning to extract degradation characteristics. Other approaches such as InstructIR (Conde et al., 2025) adopted textual guidance, and DCPT (Hu et al., 2025) proposed a novel pre-training strategy for AiOR. DFPIR (Tian et al., 2025) proposed a feature perturbation strategy for AiOR. Recent AiOR methods have adopted diffusion models. Pixel-space diffusion models (PSDMs) such as DA-CLIP (Luo et al., 2023) and DiffUIR (Zheng et al., 2024) demonstrated improved AiOR performance but lacked robust generative priors. Recent methods have utilized the strong priors of LDMs for AiOR. Diff-Plugin (Liu et al., 2024) adopts task plugins to guide an LDM for AiOR. AutoDIR (Jiang et al., 2023) automatically detects and restores degradations using an LDM. PixWizard (Lin et al., 2024) is a multi-task Lumina-next (Zhuo et al., 2024) based model capable of performing AiOR among other tasks. However, LDM-based approaches are slow at inference time, a limitation we aim to overcome using visual autoregressive modeling (VAR).
@@ 2.2 Autoregressive models in vision
Recent works (Van Den Oord et al., 2016; Tian et al., 2024) have extended autoregressive (AR) models to vision and can be categorized as pixel‐space AR (Van Den Oord et al., 2016; Van den Oord et al., 2016; Chen et al., 2018b), token‐based AR (Van Den Oord et al., 2017; Yu et al., 2023; Ramesh et al., 2021) and scale‐space AR (Tian et al., 2024; Ren et al., 2024; Guo et al., 2025). Pixel-space AR predicts raw pixels one by one in raster order, as in PixelRNN (Van Den Oord et al., 2016) and PixelCNN++ (Salimans et al., 2017), but is very slow at high resolutions. Token-based AR compresses images into discrete latent codes via vector quantization (e.g., VQ-VAE (Van Den Oord et al., 2017), VQGAN (Esser et al., 2021)) and then models code sequences with transformers (e.g. ImageGPT (Chen et al., 2020)). This trades-off codebook size and transformer capacity against tractability for high-resolution generation. Scale-space AR, as introduced in VAR (Tian et al., 2024), generates latents from coarse to fine scales and matches the quality of Diffusion Transformers (Peebles and Xie, 2023) at a fraction of the inference cost. HART (Tang et al., 2024) scales VAR to higher resolution and uses a MLP-based diffusion refiner to convert discrete VAR latents into continuous representations. Despite VAR’s success in generative tasks, it remains underexplored for image restoration with only two prior works-VarSR (Qu et al., 2025) and Varformer (Wang and Zhao, 2024). VarSR addressed super-resolution, while Varformer used VAR’s features to guide a non-generative AiOR model. In contrast, RestoreVAR is a generative model which directly trains VAR for AiOR.
@@ 3.1 Preliminaries: Visual Autoregressive Modelling
Visual Autoregressive Modelling, or VAR, is a novel autoregressive class-conditioned image generation method which uses a GPT-2 (Radford et al., 2019a) style decoder-only transformer architecture for next-scale prediction. The VAR transformer operates in the latent space of a multi-scale VQVAE which uses $K$ scales.
Degraded GT++coarse GT++fine GT
Haze
Refer to caption Refer to caption Refer to caption Refer to caption
Snow
Refer to caption Refer to caption Refer to caption Refer to caption
Rain
Refer to caption Refer to caption Refer to caption Refer to caption
Low-light
Refer to caption Refer to caption Refer to caption Refer to caption
Blur
Refer to caption Refer to caption Refer to caption Refer to caption
Given an image $I\in\mathbb{R}^{H\times W\times 3}$, the VQVAE encoder outputs a latent representation $f_{\text{cont}}\in\mathbb{R}^{H_{K}\times W_{K}\times C}$. Hereafter, we will refer to $f_{\text{cont}}$ as the continuous latent, and the latent obtained after quantization as discrete latent. Instead of directly quantizing $f_{\text{cont}}$, a multi-scale residual quantization using a shared codebook across $K$ spatial scales is performed. First, the residual and accumulated quantized (or discrete) reconstruction of $f_{\text{cont}}$ are initialized as $f_{\text{res}}^{(0)}:=f_{\text{cont}}$ and $f_{\text{quant}}^{(0)}:=0$, respectively. At each scale $k=1,\ldots,K$, an index map $r_{k}\in\mathbb{Z}^{H_{k}\times W_{k}}$ is obtained by quantizing the downsampled residual feature:
@@ 3.3.2 Detail-Preserving Restoration
$\mathcal{L}_{\text{dec}}=\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{percep}}+\lambda_{4}\mathcal{L}_{\text{adv}},\vskip-2.0pt$
@@
$\hat{f}_{\text{cont}}=f^{\text{pred}}_{\text{quant}}+\text{LRM}(f^{\text{pred}}_{\text{quant}},z),$
@@ 4.3 Comparisons
We compare RestoreVAR with state-of-the-art generative and non-generative methods for AiOR. For non-generative approaches, we include PromptIR (Potlapalli et al., 2024), InstructIR (Conde et al., 2025), AWRaCLe (Rajagopalan and Patel, 2024), DCPT (Hu et al., 2025) and DFPIR (Tian et al., 2025). Among generative methods, we compare with the LDM-based approaches Diff-Plugin (Liu et al., 2024), AutoDIR (Jiang et al., 2023) and PixWizard (Lin et al., 2024). To ensure a fair comparison, we retrained PromptIR and AWRaCLe, as their official checkpoints were not trained for most of our AiOR tasks. All other methods were evaluated using their publicly released checkpoints. For AutoDIR, we report results without the structure correction module, as this module functions as an independent, non-generative restoration network (more details in supplementary). The results reported for PixWizard were obtained using its publicly released checkpoint. We do not compare with task-specific restoration models, as RestoreVAR is proposed for the AiOR setting.
Table 4.2 presents PSNR, SSIM and LPIPS scores on the RESIDE, Snow100k, Rain13K, LOLv1 and GoPro datasets. RestoreVAR surpasses LDM-based AiOR methods at a fraction of their computational cost (inference time (s) per image)—Diff-Plugin:$2.04$s, AutoDIR: $8.477$s, PixWizard: $8.247$s and RestoreVAR: $0.201$s, highlighting the efficacy of our framework. More detailed complexity comparisons are given in the supplementary along with a derivation showing that the time complexity of VAR with maximum latent resolution $n\times n$ is $\mathcal{O}(\log n)$ lower than an LDM operating at the same latent resolution. Qualitative comparisons with LDM-based methods in Fig. 5 further illustrate that RestoreVAR produces restored images of high quality while better preserving fine details. Visual results for the Snow100k and LOLv1 datasets are provided in the supplementary. While non-generative methods achieve better scores, it is important to recognize that the performance of RestoreVAR is inherently influenced by the quality of the VAE decoder; a limitation shared by all latent generative approaches. Despite this constraint, RestoreVAR narrows the gap with non-generative methods while maintaining the benefits of a generative framework, i.e., perceptually realistic results and strong generalization capabilities. To demonstrate these strengths, we evaluate generalization using no-reference image quality metrics (following prior works (Liu et al., 2024; Jiang et al., 2023; Rajagopalan and Patel, 2024)), and assess perceptual realism through a user study.
Table 2: Quantitative comparisons of RestoreVAR against state-of-the-art non-generative and generative approaches on real-world, unseen and mixed degradations. The best result among RestoreVAR and non-generative approaches is indicated in bold.
@@
For testing generalization, we report MUSIQ (Ke et al., 2021) and CLIPIQA (Wang et al., 2023) scores in Table 2 on the real-world, unseen and mixed degradation datasets discussed in Sec. 4.2. RestoreVAR achieves higher scores than non-generative models (on average), indicating better robustness under these degradations. Qualitative results for this experiment are shown in Fig. 6, where RestoreVAR consistently outperforms non-generative approaches. Due to space constraints, qualitative comparisons with PromptIR and visual results for LOLBlur are given in the supplementary. Furthermore, RestoreVAR achieves performance competitive to those of LDM-based methods on these datasets. The slightly lower metrics for RestoreVAR can be explained by the fact that the latent diffusion backbones (Stable Diffusion (Rombach et al., 2022)) used by the competing generative models are trained on substantially larger datasets, typically hundreds of millions of images, whereas the VAR backbone is trained on only $\sim 1$M ImageNet-1K (Deng et al., 2009) images. Despite this, RestoreVAR demonstrates comparable generalization performance while offering much faster inference and improved pixel-level fidelity. To further evaluate perceptual quality, we conducted a user study in which participants rated outputs from non-generative models, AutoDIR (LDM-based) and RestoreVAR, for $50$ real-world scenes. We received $36$ responses with each participant scoring outputs based on scene consistency, restoration quality, and overall appeal on a 5-point scale. Table 3 shows that RestoreVAR received the highest average ratings (across all three criteria), highlighting its ability to produce images that align closely with human preferences. More quantitative results are provided in the supplementary.
To summarize, although non-generative AiOR models offer fast inference and good pixel-level fidelity, their generalization to real-world, unseen, and mixed degradations is limited. They work well only when test samples come from a distribution close to their training data, which hinders real-world applicability. In contrast, LDM-based generative approaches exhibit substantially stronger perceptual quality and generalization, but they are computationally expensive and suffer from low pixel-level fidelity. RestoreVAR addresses the limitations of LDM-based approaches while retaining their strengths. It achieves far superior pixel-level fidelity as established by the multi-task restoration performance on datasets in Table 4.2. Furthermore, on the real-world and unseen degradation datasets (Table 2), RestoreVAR achieves perceptual quality on par with diffusion-based methods while significantly outperforming non-generative models, demonstrating strong generalization. Moreover, RestoreVAR is substantially faster than LDM-based approaches and even nears the efficiency of some non-generative methods (see supplementary for details), despite requiring multiple steps for inference. Together, these results demonstrate that RestoreVAR closes the fidelity gap between LDM-based generative models and non-generative approaches while retaining strong perceptual quality and generalization, and offering enhanced computational efficiency.
We discuss limitations of RestoreVAR in the supplementary.
4.4 Ablations and Analysis